OpenStudy (anonymous):

What is the slope of the line whose equation is 3x + 6y = 9.

OpenStudy (anonymous):

slope of \[ax+by=c\] is \[m=-\frac{a}{b}\]

OpenStudy (anonymous):

It would not be 1/2 but instead -1/2

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

\[3x + 6y = 9 \] \[Slope = m = -\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

\[3x+6y = 9\] \[6y = -3x + 9\] \[y = \frac{ -3 }{ 6 }x + 3\] In the slope-intercpet form: y = mx + b, m is the slope. So: \[m = \frac{ -3 }{ 6 } = \frac{ -1 }{ 2 }\]